OpenStudy (anonymous):

The length of a rectangle is 4times its width. Find the area of the rectangle if the length is 12cm?

OpenStudy (anonymous):

Have you managed to figure out the width?

Parth (parthkohli):

So, the width would be 12 divided by 4. Area of a rectangle = length * width

OpenStudy (anonymous):

L = 4W Given : L = 12, then W = 12/4 = 3 Area = L*W, you have W and you have L, put in the formula above you will have your answer...